OpenStudy (anonymous):

Match each example with the appropriate property. 1. Additive Inverse 2. Additive Identity 3. Multiplicative Inverse 4. Multiplicative Identity a. (5)(1/5) = 1 b. 19 + 0 = 19 c. 4.2(1) = 4.2 d. 8 + -8 = 0

OpenStudy (midhun.madhu1987):

a * 1/a = a -----? Multiplicative inverse a+ 0 = 0 + a = a -----> Additive Identity a + (-a) = 0 -----> Additive Inverse a * 1 = a -------? Multiplicative identity